OpenStudy (anonymous):

Find the sum. (14x2 + 7x) + 9x2 =

Parth (parthkohli):

Remove the brackets. They're of no use. \( \color{Black}{\Rightarrow 14x^2 + 7x + 9x^2 }\) Now just operate on the like terms.

OpenStudy (anonymous):

14x^2+7x+9x^2 = 23x^2+7x = 23x=7=x=7/23

OpenStudy (anonymous):

sorry plz add minus sign